A generator has a 170-turn square coil, 4cm on a side.


A. Find the peak emf in this coil when it's rotating at 95.0Hz in Earth's field, B=5E-5

B=5
Answer #1

EMF=BANW


=5*10^-5*(0.04)^2*170*2*pi*95=8.117 mV


B.=6.1*(0.04)^2*170*2*pi*95=990.38 V